### v1

# Shock waves in nonlinear transmission lines

Abstract
We consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate for the ”sound” wave the coefficient of reflection from (the coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the wave speeds and the wave impedances. When only the capacitors or only the inductors are nonlinear, the coefficients are expressed in terms of the wave speeds only. We explicitly include into consideration of the shocks the dissipation, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to describe the shocks as physical objects of finite width and study their profiles. In some particular cases the profiles were obtained in terms of elementary functions.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . A very interesting particular type of signals which can propagate along such lines - the shock waves - is attracting interest since long ago landauer1 ; landauer2 . We published a series of papers on the travelling waves in Josephson transmission lines (JTL): the kinks, the solitons and the shocks kogan4 (see the last publication of that series and the references therein). The cited publication is especially relevant for what will be presented below. In that publication we considered among other problems, the small amplitude wave (”sound”) reflection from and transmission through the shock wave.
Note that from the point of view of an electrical engineer, Josephson junctions are just the nonlinear inductors, and this is how they were described in all our papers on JTL. On the other hand, the capacitors in all these papers were assumed to be linear. On the other hand, very recently we published the paper on the travelling waves in the transmission line with the linear inductors but with the nonlinear capacitors kogan5 .
In the present paper we consider the travelling waves in the transmission line where, in the general case, both the inductors and the capacitors are nonlinear. In the first part of the paper, generalizing our previous results kogan4 ; kogan5 , we calculate the ”sound” reflection from and transmission through the shock wave in these nonlinear transmission lines. In the second part of the paper we include explicitly the ohmic losses (which are of course necessary for the existence of the shock waves) into the circuit equations. This allows us to study the profile of the shock wave.
II The circuit equations
The transmission line constructed from the identical nonlinear inductors and the identical nonlinear capacitors is shown on Fig. 1.
Figure 1: Lossless nonlinear transmission line
We take the voltages on the capacitors $V_{n}$ and the current through the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
$\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1},$ (1a)
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
where
[…]
$Q_{n}$ is the charge of the nonlinear capacitor and ${\cal V}_{n}$ is the voltage on the nonlinear inductor.
In the continuum approximation we treat $n$ as the continuous variable $x$ and approximate the finite differences in the r.h.s. of the equations by the first derivatives with respect to $x$, after which the equations take the form
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x},$ (3a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}$ (3b)
(we measure distance in the units of the transmission line period). To close the system (3) we should specify
$Q$ $=Q(V),$ (4a)
$\Phi$ $=\Phi(I).$ (4b)
Further on we’ll consider $Q$ and $\Phi$ as known functions of $V$ and $I$ respectively. Note however, that sometimes it will be convenient for us to invert the relations (4) and consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively. We’ll switch between the representations without further notice.
III The ”sound” and the shocks
Because (3) is nonlinear, it can have a lot of different types of solutions. In this paper we’ll be interested in only two types of those. First type - small amplitude waves on a homogeneous background $\varphi_{0},V_{0}$.
$V$ $=V_{0}+v,$ (5a)
$I$ $=I_{0}+i.$ (5b)
For such waves Eq. (3) is simplified to
$C_{0}\frac{\partial v}{\partial t}$ $=-\frac{\partial i}{\partial x},$ (6a)
$L_{0}\frac{\partial i}{\partial t}$ $=-\frac{\partial v}{\partial x},$ (6b)
where
$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{V=V_{0}},$ (7a)
$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{I=I_{0}}.$ (7b)
The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$, where
$u_{0}^{2}=\frac{1}{L_{0}C_{0}}$ (8)
is the ”sound” speed. The current and the voltage in the waves are connected by the equations
$v=\pm Z_{0}i,$ (9)
where
$Z_{0}=\sqrt{\frac{L_{0}}{C_{0}}}$ (10)
is the linear impedance of the transmission line.
The second type of solutions, we are interested in, is shock waves, satisfying the conditions (see Section V)
$U_{21}\left(Q_{2}-Q_{1}\right)$ $=I_{2}-I_{1},$ (11a)
$U_{21}\left(\Phi_{2}-\Phi_{1}\right)$ $=V_{2}-V_{1},$ (11b)
where the indices 1 and 2 refer to the quantities before and after the shock respectively, and $U_{21}$ is the shock wave speed. Equations (11) mean, in particular, that from the five parameters ($V_{1},V_{2},I_{1},I_{2},U$) - three are the free parameters, and the other two are defined by the equations.
The indices 1 and 2 enter in a symmetrical way in the system 11, so Eqs. 11 as they stand can describe propagation of the shock in both directions. However, in Section V) we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
$u_{2}^{2}>U_{21}^{2}>u_{1}^{2}.$ (12)
IV The ”sound” reflection and transmission
IV.1 Nonlinear transmission line
We’ll be interested in two problems landau ; congy . The first one: A ”sound” wave is incident from the rear on a shock wave (see Fig. 2).
Figure 2: Reflection of a ”sound” wave from a shock wave.
The second problem: A ”sound” wave is incident from the front on a shock wave (see Fig. 3).
Figure 3: Transmission of a ”sound” wave through a shock wave.
(Note the difference between the problems we consider and the problem of propagation of high-frequency wave packets along a smooth evolving background flow kamchatnov .) While formulating both problems we took into account the inequalities (12), meaning that there is no ”sound” transmission in the first case, and no ”sound” reflection in the second case.
For the first problem mentioned above we have
$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}&V_{a}+v^{(in)}+v^{(r)}\\ I_{b}&I_{a}+i^{(in)}+i^{(r)}\end{array}\right),$ (17)
where (in) stands for the incident ”sound” wave and (r) for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
$U_{21}=U_{ab}+\delta U^{(refl)}.$ (18)
Substituting (17) into (11), in linear with respect to the ”sound” wave amplitudes approximation we obtain
$\delta U^{(refl)}\left(Q_{a}-Q_{b}\right)+U_{ab}C_{a}\left(v^{(in)}+v^{(r)}\right)=i^{(in)}+i^{(r)},$ (19a)
$\delta U^{(refl)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}L_{a}\left(i^{(in)}+i^{(r)}\right)=v^{(in)}+v^{(r)}.$ (19b)
Taking into account the relations
$v^{(in)}$ $=Z_{a}i^{(in)},$ (20a)
$v^{(r)}$ $=-Z_{a}i^{(r)}$ (20b)
(the difference in the signs is because of the opposite directions of propagation of the two waves) and dividing the equations (19) to each other we obtain
$Z_{ab}=Z_{a}\frac{u_{in}i^{(in)}+\left(u_{a}+U_{ab}\right)i^{(r)}}{u_{r}i^{(in)}-\left(u_{a}+U_{ab}\right)i^{(r)}},$ (21)
where $u_{in}=u_{a}-U$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
Solving (21) we obtain for the reflection coefficient
$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{u_{in}\left(Z_{ab}-Z_{a}\right)}{u_{r}\left(Z_{ab}+Z_{a}\right)}.$ (23)
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
For the second problem we have
$\left(\begin{array}[]{c|c}V_{1}&V_{2}\\ I_{1}&I_{2}\end{array}\right)=\left(\begin{array}[]{c|c}V_{b}+V^{(in)}&V_{a}+v^{(t)}\\ I_{b}+i^{(in)}&I_{a}+i^{(t)}\end{array}\right),$ (28)
where (t) stands for the transmitted wave. Substituting (28) into (11), in linear approximation we obtain
$\delta U^{(trans)}\left(Q_{a}-Q_{b}\right)+U_{ab}\left(C_{a}v^{(t)}-C_{b}v^{(in)}\right)=i^{(t)}-i^{(in)},$ (29a)
$\delta U^{(trans)}\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(L_{a}i^{(t)}-L_{b}i^{(in)}\right)=v^{(t)}-v^{(in)}.$ (29b)
Taking into account the relations
$v^{(in)}$ $=-Z_{b}i^{(in)},$ (30a)
$v^{(t)}$ $=-Z_{a}i^{(t)}$ (30b)
and dividing the equations (29) to each other we obtain
$Z_{ab}=\frac{u_{a}u_{in}Z_{b}i^{(in)}-u_{b}u_{t}Z_{a}i^{(t)}}{-u_{a}u_{in}i^{(in)}+u_{b}u_{t}i^{(t)}},$ (31)
where $u_{in}=u_{b}+U$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U$ is the speed of the transmitted ”sound” wave relative to the shock wave. Solving (31) we obtain for the transmission coefficient
$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)}{u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)}.$ (32)
As one could have expected, the transmission coefficient goes to 1 when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
Note that Eqs. (21) and (31) can be written down respectively in the form
$u_{in}\left(Z_{ab}-Z_{a}\right)i^{(in)}$ $=u_{r}\left(Z_{ab}+Z_{a}\right)i^{(r)},$ (33a)
$u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)i^{(in)}$ $=u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)i^{(t)}.$ (33b)
We wonder whether these equations could have been written down at an earlier stage by considering propagation of the waves in the frame of reference moving with the shock wave.
IV.2 Half-nonlinear transmission line
IV.2.1 Linear capacitor
For linear capacitor ($Q=VC$, where $C$ is a constant) from (8) and (10) follows that
$Z_{a,b}=\frac{1}{Cu_{a,b}},$ (34)
and from (11) and (22) follows that
$Z_{ab}=\frac{1}{CU_{ab}}.$ (35)
Hence Eqs. (23) and (32) can be simplified respectively to
$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (36a)
$T$ $=\frac{u_{a}^{2}u_{in}^{2}}{u_{b}^{2}u_{t}^{2}},$ (36b)
and from (8) and (11) follows
$u_{a,b}^{2}$ $=\frac{1}{C}\left.\frac{dI}{d\Phi}\right|_{\Phi=\Phi_{a,b}}$ (37a)
$U_{ab}^{2}$ $=\frac{1}{C}\frac{I_{a}-I_{b}}{\Phi_{a}-\Phi_{b}}.$ (37b)
Consider the case of JTL kogan4 . In this case from the second Josephson law
${\cal V}=\frac{\hbar}{2e}\frac{d\varphi}{dt},$ (38)
where $\varphi$ is Josephson phase, we obtain
$\Phi=\frac{\hbar}{2e}\varphi.$ (39)
Hence the first Josephson law gives
$I=I_{c}\sin\left(\frac{2e\Phi}{\hbar}\right),$ (40)
and we can easily calculate all the velocities. Equations (36) for the case of JTL were obtained in our previous publication kogan4 . (Comparing the results of Ref. kogan4 with those presented in this paper, one should pay attention to a slightly different definitions of the transmission and the reflection coefficients in the two papers.)
IV.2.2 Linear inductor
For linear inductor ($\Phi=LI$, where $L=$ is a constant) from (8) and (10) follows that
$Z_{a,b}=Lu_{a,b},$ (41)
and from (11) and (22) follows that
$Z_{ab}=LU_{ab}.$ (42)
Hence Eqs. (23) and (32) can be respectively simplified to
$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (43a)
$T$ $=\frac{u_{a}u_{in}^{2}}{u_{b}u_{t}^{2}},$ (43b)
and from (8) and (11) follows
$u_{a,b}^{2}$ $=\frac{1}{L}\left.\frac{dV}{dQ}\right|_{Q=Q_{a,b}}$ (44a)
$U_{ab}^{2}$ $=\frac{1}{L}\frac{V_{a}-V_{b}}{Q_{a}-Q_{b}}.$ (44b)
Note that because of the symmetry of Eq. (3) with respect to simultaneous interchange
$Q$ $\leftrightarrows\Phi,$ (45a)
$V$ $\leftrightarrows I,$ (45b)
Eqs. (43), (44) follow from Eqs. (36), (37).
V Lossy transmission line
V.1 Nonlinear transmission line
As it is well known landau , the shock waves exist only in the presence of dissipation. (In the absence of the dissipation, the localized travelling waves in the systems we consider are kinks and solitons kogan1 ; kogan2 ; kogan4 .) On the other hand, many properties of the shocks in fluids were studied beginning from the XIX century without taking the dissipation explicitly into account, but just postulating the existence of discontinuous solutions and the jump conditions krehl . This is exactly what we did in the previous Sections.
However in this Section, to be consistent, we assume the existence of ohmic resistors in the transmission line, shunting the inductors and in series with the capacitors. Such transmission line is shown on Fig. 4. We’ll resolve the zero width discontinuities from the previous Sections, which will turn out to be narrow (for weak dissipation) crossover regions and rigorously prove the jump conditions.
Figure 4: Lossy nonlinear transmission line
The corrected Eqs. (1) are
$\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1}+\frac{1}{R_{L}}\frac{d}{dt}\left(\Phi_{n}-\Phi_{n+1}\right),$ (46a)
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1}+R_{C}\frac{d}{dt}\left(Q_{n}-Q_{n+1}\right).$ (46b)
In the continuum approximation (46) takes the form (compare with (3))
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x}-\frac{1}{R_{L}}\frac{\partial^{2}\Phi}{\partial x\partial t},$ (47a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}-R_{C}\frac{\partial^{2}Q}{\partial x\partial t}.$ (47b)
Further on we’ll consider the travelling waves, for which all the dependent variables depend upon the single independent variable $\tau=t-x/U$, where $U$ is the speed of the wave. For such waves (47) turns into the system of ODE which after integration takes the form
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const},$ (48a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const},$ (48b)
To study the shock waves we should impose the boundary conditions
$\lim_{\tau\to-\infty}Q(\tau)$ $=Q_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}Q(\tau)=Q_{2};$ (49a)
$\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}\Phi(\tau)=\Phi_{2}.$ (49b)
Both $(Q_{1},\Phi_{1})$ and $(Q_{2},\Phi_{2})$ are the fixed points (in the $(Q,\Phi)$ phase space), hence the r.h.s. of (48b) is equal to zero in both points. Equating these two values we recover the jump conditions (11b). Doing the same procedure with (48a) we recover the jump conditions (11a).
The point $(Q_{1},\Phi_{1})$ is the unstable fixed point of the system (48). This means
$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (52)
or
$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{1}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{1}}\end{array}\right|<0.$ (56)
Calculating the determinant and recalling the result for the ”sound” speed we obtain
$U_{21}^{2}>u_{1}^{2}.$ (57)
On the other hand, the point $(Q_{2},\Phi_{2})$ is the stable fixed point, hence we obtain
$u_{2}^{2}>U_{21}^{2}.$ (58)
Thus we proved the inequalities (12) postulated previously.
V.2 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix A) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (48) given by (45) (to which we should add $1/R_{L}\leftrightarrows R_{C}$), the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
Because of the relation $\Phi=LI$, where $L$ is a constant, we can exclude the current from (48) and after a bit of algebra obtain
$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV}{d\tau}=F,$ (59)
where
$F\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (60)
Equation (59) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (60). The boundary conditions (49a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$, thus it is crucial to keep the nonlinear dependence of $Q$ upon $V$ in 60. We may present these maximum/minimum conditions as
$\left.\frac{dF(Q)}{dQ}\right|_{Q=Q_{2}}=\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}-\left.\frac{dV}{dQ}\right|_{Q=Q_{2}}<0$ (61)
and opposite inequality for $V=V_{1}$. Recalling (44) we realise that we have reproduced in this particular case Eqs. (57), (58).
Equation (59) is simplified (in a trivial way) when there is a leading mechanism of dissipation, that is when either $R_{C}Q\ll LV/R_{L}$ or $R_{C}Q\gg LV/R_{L}$. In each of these cases, in the l.h.s. of (59) a single term with the first derivative should be left, and the equation can be easily integrated in quadratures (in some cases in elementary functions). Thus we obtained Eqs. (48a) and (48b) of our paper kogan3 .
V.2.1 Weak shocks
Alternatively, Eq. (59) is simplified (in a less trivial way) for weak shock, when $|V_{1}-V_{2}|$ is small. We can put in the l.h.s. of the equation
$\frac{dQ}{d\tau}=C_{s}\frac{dV}{d\tau},$ (62)
where $C_{s}$ is a differential capacitance assumed to be constant for the voltages considered, and (59) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=F(V),$ (63)
where $\tau_{R}\equiv R_{C}C_{s}+L/R_{L}$ and $T^{2}\equiv LC_{s}R_{C}/R_{L}$. The solution of (63) should satisfy the boundary conditions (49a), rewritten as
$\lim_{\tau\to-\infty}V(\tau)=V_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}V(\tau)=V_{2}.$ (64)
Before going further it is convenient to rewrite (60) as
$F(V)=\frac{(V-V_{1})\left[Q(V)-Q_{2}\right]-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (65)
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks we can approximate $Q(V)$ in (65) as
$Q(V)=Q_{0}+C_{0}(V+\beta V^{2}),$ (66)
and (63) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\beta(V-V_{1})(V-V_{2}).$ (67)
In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(V_{1}-V_{2})}{6},$ (68)
there exists the elementary solution of (67)
$V=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (69)
where
$\chi=\frac{\tau_{R}}{5T^{2}}.$ (70)
Approximation (66) can be inadequate, if $V_{1},V_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (66) to
$Q=Q_{0}+C_{0}(V+\beta V^{2}+\gamma V^{3}).$ (71)
In this case (63) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\gamma(V-V_{1})(V-V_{2})(V+V_{3}),$ (72)
where $V_{3}=\beta/\gamma+V_{1}+V_{2}$.
In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(V_{1}+V_{2}+2V_{3}\right)^{2}}{2},$ (73)
there exists the elementary solution of (72)
$V=V_{2}+\frac{V_{1}-V_{2}}{\exp\left(\psi\tau\right)+1},$ (74)
where
$\psi=\frac{\tau_{R}\left(V_{1}-V_{2}\right)}{T^{2}\left(V_{1}+V_{2}+2V_{3}\right)}.$ (75)
Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A Shock profile for the fully nonlinear transmission line
Taking into account the boundary conditions we can present (48) as
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)-UQ_{1}+I_{1},$ (76a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)-U\Phi_{1}+V_{1},$ (76b)
Equation (76) allows us to study the profile of the shock wave in the general case.
Equation (76) is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (76) either $R_{C}=0$ or $R_{L}=\infty$. In the first case we equate the r.h.s. of (76b) to zero and obtain
$V(Q)=U\Phi-U\Phi_{1}+V_{1}.$ (77)
Solving (77) with respect to $Q$ and substituting into (76a) we obtain the equation which can be easily integrated in quadratures.
In the second case we equate the r.h.s. of (76a) to zero and obtain
$I(\Phi)=UQ-UQ_{1}+I_{1}.$ (78)
Solving (78) with respect to $\Phi$ and substituting into (76b) we obtain the equation which can be integrated in quadratures.
If both $R_{C}>0$ and $R_{L}<\infty$, from (76) we may obtain the equation
$\frac{1}{R_{L}R_{C}}\frac{d\Phi}{dQ}=\frac{UQ-I(\Phi)-UQ_{1}+I_{1}}{U\Phi-V(Q)-U\Phi_{1}+V_{1}}.$ (79)
If we can solve (79), we can substitute the solution into one of the equations (76) and integrate the resulting equations in quadratures to obtain the profile of the shock wave.
Appendix B Integration of the ODE which don’t contain explicitly the independent variable in terms of elementary functions
Looking at Eq. (63) we realize that it doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $V$ as the new independent variable and $E\equiv dV/d\tau$ as the new dependent variable. In the new variables (63) takes the form
$T^{2}E\frac{dE}{dV}+\tau_{R}E=F(V).$ (80)
As it follows from (49a), the boundary conditions for (80) are
$\lim_{V\to V_{1},V_{2}}E(V)=0.$ (81)
Note that Eq. (80) and the boundary conditions (81) are symmetric with respect to interchange of the indexes $2$ and $1$, while Eq. (63) and the boundary conditions (49a) - certainly not. So to the boundary conditions (64) we should add the demand (61) (or opposite inequality for $V=V_{1}$). Substituting (61) into (80) we obtain the demand written in the new variables as
$\left.\frac{T^{2}}{2}\frac{d^{2}E}{dV^{2}}\right|_{V=V_{2}}+\left.\tau_{R}\frac{dE}{dV}\right|_{V=V_{2}}<0.$ (82)
Since (80) is the first order differential equation, the problem is overdetermined. However, for the problems considered explicitly in Section V.2.1, not only will we be able to find the solutions of (80) satisfying the boundary conditions (81) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
For $Q(V)$ given by (71), Eq. (80) can be presented as
$T^{2}E\frac{dE}{dV}+\tau_{R}E=\gamma(V-V_{1})(V-V_{2})(V+V_{3}).$ (83)
The boundary conditions (81) and the r.h.s. of (83) prompt the idea to look for $E$ as the polynomial function proportional to $\left(V-V_{1}\right)\left(V-V_{2}\right)$. Because the r.h.s. of (83) is the third order polynomial of $V$, the multiplying factor should be a constant. That is the natural inspired guess is kogan4 ; kogan5
$E(V)=\frac{\psi}{V_{1}-V_{2}}\left(V-V_{1}\right)\left(V-V_{2}\right)$ (84)
(the normalization was chosen for later convenience).
Substituting (84) into (83) we obtain
$\frac{T^{2}\psi^{2}}{\left(V_{1}-V_{2}\right)^{2}}\left(2V-V_{1}-V_{2}\right)+\frac{\tau_{R}\psi}{V_{1}-V_{2}}=\gamma\left(V+V_{3}\right).$
Hence the condition on the parameters of the problem, for (84) to be the solution of (83) is given by (73), and the constant multiplier in (84) should is given by (75). Substituting (84) into the equation
$\frac{dV}{d\tau}=E(V)$ (86)
we obtain the solution (74).
For $Q(V)$ given by (66) we obtain, instead of (83), Abel equation of the second kind polyanin
$T^{2}E\frac{dE}{dV}+\tau_{R}E=\beta(V-V_{1})(V-V_{2}).$ (87)
Counting the powers of $V$ in the r.h.s. of (87) we come to the conclusion that if $E$ is a polynomial function of $V$, it should be a 3/2 order polynomial function. And the boundary conditions (81) tell us that this polynomial should zero out at $V=V_{1,2}$. To acquire an inspired guess how the polynomial may look like, let us return to the solution of (83) we obtained previously and split in two one multiplier in the solution
$\frac{V-V_{1}}{V_{1}-V_{2}}=\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}+1\right].$
All that was said above being taken together, prompts us to check up the solution of (87) in the form kogan4 ; kogan5
$E(V)=2\chi\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left(V-V_{2}\right).$ (89)
Substituting the ansatz into (87) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (68) and $\chi$ is given by (70). Substituting (89) into (86) we obtain (69)
Actually, because of the symmetry of Eq. (80) and the boundary conditions (81) with respect to interchange of the indexes $2$ and $1$, instead of (89) and the condition (68) we could as well have obtained those with the indexes $2$ and $1$ interchanged. However, one can easily see that (89) satisfies the demand (82), and (89) with the indexes 1 and 2 interchanged - doesn’t.