### v4

# Shock waves in nonlinear transmission lines

Abstract
In the first half of the paper we consider interaction between the small amplitude travelling waves (”sound”) and the shock waves in the transmission line containing both nonlinear capacitors and nonlinear inductors. We calculate the ”sound” wave coefficient of reflection from (coefficient of transmission through) the shock wave. These coefficients are expressed in terms of the speeds of the ”sound” waves relative to the shock and the wave impedances. In the second half of the paper we explicitly include into consideration the dissipation in the system, introducing ohmic resistors shunting the inductors and also in series with the capacitors. This allows us to justify the conditions on the shocks, postulated in the first half of the paper. This also allows us to describe the shocks as physical objects of finite width and study their profiles. In some particular cases the profiles were obtained in terms of elementary functions.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . A very interesting particular type of signals which can propagate along such lines - the shock waves - is attracting interest since long ago landauer1 ; landauer2 . We published a series of papers on the travelling waves in Josephson transmission lines (JTL): the kinks, the solitons and the shocks kogan4 (see the last publication of that series and the references therein). The cited publication is especially relevant for what will be presented below. In that publication we considered among other problems, the small amplitude wave (”sound”) reflection from and transmission through the shock wave.
Note that from the point of view of an electrical engineer, Josephson junctions are just the nonlinear inductors, and this is how they were described in all our papers on JTL. On the other hand, the capacitors in all these papers were assumed to be linear. On the other hand, very recently we published the paper on the travelling waves in the transmission line with the linear inductors but with the nonlinear capacitors kogan5 .
In the present paper we consider the travelling waves in the transmission line where, in the general case, both the inductors and the capacitors are nonlinear. In the first part of the paper, generalizing our previous results kogan4 ; kogan5 , we calculate the ”sound” reflection from and transmission through the shock wave in these nonlinear transmission lines. In the second part of the paper we include explicitly the ohmic losses (which are of course necessary for the existence of the shock waves) into the circuit equations.
II The circuit equations
The transmission line constructed from the identical nonlinear inductors and the identical nonlinear capacitors is shown on Fig. 1.
Figure 1: Lossless nonlinear transmission line
We take the capacitors charges $Q_{n}$ and the integrated voltages on the inductors
[…]
as the dynamical variables. The circuit equations (Kirchhoff laws) are
$\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1},$ (2a)
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1}.$ (2b)
To close the system (2) we should specify the connection between the voltages on the capacitors $V_{n}$ and the charges and also between the currents through the inductors $I_{n}$ and the integrated voltages
$V_{n}$ $=V(Q_{n}),$ (3a)
$I_{n}$ $=I(\Phi_{n}).$ (3b)
Further on we’ll consider $V(Q)$ and $I(\Phi)$ as known functions.
In the continuum approximation we treat $n$ as the continuous variable $x$ (we measure distance in the units of the transmission line period) and approximate the finite differences in the r.h.s. of the equations by the first derivatives with respect to $x$, after which the equations take the form
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x},$ (4a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}.$ (4b)
III The ”sound” and the shocks
Because (4) is nonlinear, it can have a lot of different types of solutions. In this paper we’ll be interested in only two types of those. First type - small amplitude waves on a homogeneous background $Q_{0},\Phi_{0}$.
$Q$ $=Q_{0}+q,$ (5a)
$\Phi$ $=\Phi_{0}+\phi.$ (5b)
For such waves Eq. (4) is simplified to
$\frac{\partial q}{\partial t}$ $=-\frac{1}{L_{0}}\frac{\partial\phi}{\partial x},$ (6a)
$\frac{\partial\phi}{\partial t}$ $=-\frac{1}{C_{0}}\frac{\partial q}{\partial x},$ (6b)
where
[…]
The solutions of Eq. (6) (which, for brevity we’ll call the ”sound”) are right- and left-propagating travelling waves, each depending upon the single variable $\tau_{\pm}=t\mp x/u_{0}$. The variables in the ”sound” wave are connected by the equations
$v$ =±$=\pm u_{0}\phi,$ (8a)
$i$ =±$=\pm u_{0}q,$ (8b)
$\phi$ =±$=\pm Z_{0}q,$ (8c)
where
$u_{0}$ $=\frac{1}{\sqrt{L_{0}C_{0}}},$ (9a)
$Z_{0}$ $=\sqrt{\frac{L_{0}}{C_{0}}}$ (9b)
are the ”sound” speed and the linear impedance of the transmission line respectively.
The second type of solutions, we are interested in, is shock waves, containing (in the approximation of the present Section) discontinuities, satisfying the conditions (see Section V)
$U_{21}\left(Q_{2}-Q_{1}\right)$ $=I_{2}-I_{1},$ (10a)
$U_{21}\left(\Phi_{2}-\Phi_{1}\right)$ $=V_{2}-V_{1},$ (10b)
where $V_{a,b}\equiv V(Q_{a,b})$, $I_{a,b}\equiv I(\Phi_{a,b})$, and $U_{21}$ is the shock wave speed. The indices 1 and 2, which refer to the quantities before and after the shock respectively, enter in the system (10) in a symmetrical way. However, in Section V we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
$u_{2}^{2}>U_{21}^{2}>u_{1}^{2}.$ (11)
The shock is supersonic with respect to the phase before the front, but subsonic with respect to the phase behind the front.
IV The ”sound” reflection and transmission
We’ll be interested in two problems landau ; congy . The first one: A ”sound” wave is incident from the rear on a shock wave (see Fig. 2).
Figure 2: Reflection of a ”sound” wave from a shock wave.
The second problem: A ”sound” wave is incident from the front on a shock wave (see Fig. 3).
Figure 3: Transmission of a ”sound” wave through a shock wave.
(Note the difference between the problems we consider and the problem of propagation of high-frequency wave packets along a smooth evolving background flow kamchatnov .) While formulating both problems we took into account the inequalities (11), meaning that there is no ”sound” transmission in the first case, and no ”sound” reflection in the second case.
For the first problem mentioned above we have
$\left(\begin{array}[]{c|c}Q_{1}&Q_{2}\\ \Phi_{1}&\Phi_{2}\end{array}\right)=\left(\begin{array}[]{c|c}Q_{b}&Q_{a}+q^{(in)}+q^{(r)}\\ \Phi_{b}&\Phi_{a}+\phi^{(in)}+\phi^{(r)}\end{array}\right),$ (16)
where (in) stands for the incident and (r) - for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
$U_{21}=U_{ab}+\delta U$ (17)
(same in the next problem). Substituting (16) into (10), in linear with respect to the ”sound” wave amplitudes approximation we obtain (taking into account (8a) and (8b))
$\delta U\left(Q_{a}-Q_{b}\right)+U_{ab}\left(q^{(r)}+q^{(in)}\right)=u_{a}\left(q^{(in)}-q^{(r)}\right),$ (18a)
$\delta U\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(\phi^{(r)}+\phi^{(in)}\right)=u_{a}\left(\phi^{(in)}-\phi^{(r)}\right).$ (18b)
Excluding $\delta U$ from (18) and taking into account (8c) we obtain for the reflection coefficient
$R\equiv\frac{\phi^{(r)}}{\phi^{(in)}}=\frac{\left(Z_{a}-Z_{ab}\right)u_{in}}{\left(Z_{a}+Z_{ab}\right)u_{r}},$ (19)
where $u_{in}=u_{a}-U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U_{ab}$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
$Z_{ab}\equiv\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (20)
Note that from (10) follows
$\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}=\frac{V_{a}-V_{b}}{I_{a}-I_{b}}.$ (21)
For the second problem we have
$\left(\begin{array}[]{c|c}Q_{1}&Q_{2}\\ \Phi_{1}&\Phi_{2}\end{array}\right)=\left(\begin{array}[]{c|c}Q_{b}+q^{(in)}&Q_{a}+q^{(t)}\\ \Phi_{b}+\phi^{(in)}&\Phi_{a}+\phi^{(t)}\end{array}\right),$ (26)
where (t) stands for the transmitted wave. Substituting (26) into (10), in linear approximation we obtain
$\delta U\left(Q_{a}-Q_{b}\right)+U_{ab}\left(q^{(t)}-q^{(in)}\right)=u_{b}q^{(in)}-u_{a}q^{(t)},$ (27a)
$\delta U\left(\Phi_{a}-\Phi_{b}\right)+U_{ab}\left(\phi^{(t)}-\phi^{(in)}\right)=u_{b}\phi^{(in)}-u_{a}\phi^{(t)}.$ (27b)
Excluding $\delta U$ from (27) and taking into account (8c) we obtain for the transmission coefficient
$T\equiv\frac{\phi^{(t)}}{\phi^{(in)}}=\frac{\left(Z_{b}+Z_{ab}\right)Z_{a}u_{in}}{\left(Z_{a}+Z_{ab}\right)Z_{b}u_{t}},$ (28)
where $u_{in}=u_{b}+U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U_{ab}$ is the speed of the transmitted ”sound” wave relative to the shock wave.
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $Q_{a}\to Q_{b},\Phi_{a}\to\Phi_{b}$, in other words, when the shock wave itself is nearly the ”sound” wave. Similarly, in these circumstances the transmission coefficient goes to 1. Looking at (19) and (28) we wonder whether these equations could have been obtained by considering propagation of the waves in the frame of reference moving with the shock wave.
IV.1 Half-nonlinear transmission line
IV.1.1 Linear capacitor
For linear capacitor ($V=Q/C$, where $C$ is a constant)
$Z_{a,b}=\frac{1}{Cu_{a,b}},\hskip 28.45274ptZ_{ab}=\frac{1}{CU_{ab}}.$ (29)
Hence Eqs. (19) and (28) can be simplified respectively to
$R$ $=-\frac{u_{in}^{2}}{u_{r}^{2}},$ (30a)
$T$ $=\frac{u_{in}^{2}}{u_{t}^{2}},$ (30b)
and from (9a) and (10) follows
$u_{a,b}^{2}$ $=\frac{1}{C}\left.\frac{dI}{d\Phi}\right|_{\Phi=\Phi_{a,b}}$ (31a)
$U_{ab}^{2}$ $=\frac{1}{C}\frac{I_{a}-I_{b}}{\Phi_{a}-\Phi_{b}}.$ (31b)
For example, for the JTL
$I=I_{c}\sin\left(\frac{2e\Phi}{\hbar}\right),$ (32)
and all the velocities can be easily calculated. Equations (30) in this case were obtained in our previous publication kogan4 .
IV.1.2 Linear inductor
For linear inductor ($I=\Phi/L$, where $L$ is a constant)
$Z_{a,b}=Lu_{a,b},\hskip 28.45274ptZ_{ab}=LU_{ab}.$ (33)
Hence Eqs. (19) and (28) can be respectively simplified to
$R$ $=\frac{u_{in}^{2}}{u_{r}^{2}},$ (34a)
$T$ $=\frac{Z_{a}u_{in}^{2}}{Z_{b}u_{t}^{2}},$ (34b)
and from (9a) and (10) follows
$u_{a,b}^{2}$ $=\frac{1}{L}\left.\frac{dV}{dQ}\right|_{Q=Q_{a,b}}$ (35a)
$U_{ab}^{2}$ $=\frac{1}{L}\frac{V_{a}-V_{b}}{Q_{a}-Q_{b}}.$ (35b)
V Lossy transmission line
As it is well known landau , the shock waves exist only in the presence of dissipation. (In the absence of the dissipation, the localized travelling waves in the systems we consider are kinks and solitons kogan1 ; kogan2 ; kogan4 .) On the other hand, many properties of the shocks in fluids were studied beginning from the XIX century without taking the dissipation explicitly into account, but just postulating the existence of discontinuous solutions and the jump conditions krehl . This is exactly what we did in the previous Sections.
However in this Section, to be consistent, we assume the existence of ohmic resistors in the transmission line, shunting the inductors and in series with the capacitors. Such transmission line is shown on Fig. 4. We’ll resolve the zero width discontinuities from the previous Sections, which will turn out to be narrow (for weak dissipation) crossover regions and rigorously prove the jump conditions.
Figure 4: Lossy nonlinear transmission line
The corrected Eqs. (2) are
$\frac{dQ_{n}}{dt}$ $=I_{n}-I_{n+1}+\frac{1}{R_{L}}\frac{d}{dt}\left(\Phi_{n}-\Phi_{n+1}\right),$ (36a)
$\frac{d\Phi_{n+1}}{dt}$ $=V_{n}-V_{n+1}+R_{C}\frac{d}{dt}\left(Q_{n}-Q_{n+1}\right).$ (36b)
In the continuum approximation (36) takes the form (compare with (4))
$\frac{\partial Q}{\partial t}$ $=-\frac{\partial I}{\partial x}-\frac{1}{R_{L}}\frac{\partial^{2}\Phi}{\partial x\partial t},$ (37a)
$\frac{\partial\Phi}{\partial t}$ $=-\frac{\partial V}{\partial x}-R_{C}\frac{\partial^{2}Q}{\partial x\partial t}.$ (37b)
Further on we’ll consider the travelling waves, for which all the dependent variables depend upon the single independent variable $\tau=t-x/U$, where $U$ is the speed of the wave. For such waves (37) turns into the system of ODE which after integration takes the form
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)+\text{const},$ (38a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)+\text{const},$ (38b)
To study the shock waves we should impose the boundary conditions
$\lim_{\tau\to-\infty}Q(\tau)$ $=Q_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}Q(\tau)=Q_{2};$ (39a)
$\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}\Phi(\tau)=\Phi_{2}.$ (39b)
Equating the r.h.s. of (38a) at $\tau=\pm\infty$, we recover the jump conditions (10a). Doing the same procedure with (38b) we recover the jump conditions (10b). Because the quantities ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$), satisfy the constraint (21), Eq. (39) specifies 3 boundary conditions, which thus define also the shock speed $U$. Continuing the analysis of the boundary conditions, we notice that $(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (38), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
$\left|\begin{array}[]{cc}\frac{d}{d\Phi}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}Q-I(\Phi)\right]_{Q_{1},\Phi_{1}}\\ \frac{d}{d\Phi}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}&\frac{d}{dQ}\left[U_{21}\Phi-V(Q)\right]_{Q_{1},\Phi_{1}}\end{array}\right|<0,$ (42)
or
$\left|\begin{array}[]{cc}-\left.\frac{dI}{d\Phi}\right|_{\Phi_{1}}&U_{21}\\ U_{21}&-\left.\frac{dV}{dQ}\right|_{Q_{1}}\end{array}\right|<0.$ (46)
Calculating the determinant and recalling the result for the ”sound” speed we obtain
$U_{21}^{2}>u_{1}^{2}.$ (47)
Similarly for the point $(Q_{2},\Phi_{2})$ we obtain
$u_{2}^{2}>U_{21}^{2}.$ (48)
Thus we proved the inequalities (11) postulated previously.
Note that the kinks, which we considered in our previous publications in addition to shocks but do not consider here, correspond to both fixed points being semistable. Thus the kinks turn out to be supersonic both with respect to the phase before the front and the phase behind the front.
V.1 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (38) the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section.. We can exclude the current from (38) and after a bit of algebra obtain
$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV(Q)}{d\tau}=F(Q),$ (49)
where
$F(Q)\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V(Q)+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (50)
To (49) we should add the boundary conditions (39a). (The third arbitrary constant appears when we connect $\Phi$ with the found $Q$.)
Equation (49) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (50). The boundary conditions (39a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$. We may present these maximum/minimum conditions as
$\left.\frac{dF(Q)}{dQ}\right|_{Q=Q_{2}}=\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}-\left.\frac{dV}{dQ}\right|_{Q=Q_{2}}<0$ (51)
and opposite inequality for $V=V_{1}$. Recalling (35) we realise that we have reproduced in this particular case Eqs. (47), (48).
V.1.1 Weak shocks
Before going further it is convenient to rewrite (50) as
$F(Q)=\frac{\left[V(Q)-V_{1}\right]\left(Q-Q_{2}\right)-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (52)
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
$V(Q)=\text{const}+\frac{Q+\beta Q^{2}}{C_{s}},$ (53)
where $\beta$ and $C_{s}$ are assumed to be constant. We also may take $V(Q)$ in the last term in the l.h.s. of (49) in the linear approximation. So finally Eq. (49) becomes
$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\beta(Q-Q_{1})(Q-Q_{2}),$ (54)
where
$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}},$ (55a)
$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}}.$ (55b)
One can check up that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(Q_{1}-Q_{2})}{6},$ (56)
there exists the elementary solution of (54)
$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (57)
where
$\chi=\frac{\tau_{R}}{5T^{2}}.$ (58)
Approximation (53) can be inadequate, if $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (53) to
$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}}.$ (59)
In this case (49) becomes
$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}),$ (60)
where $Q_{3}=\beta/\gamma+Q_{1}+Q_{2}$. One can check up that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(Q_{1}+Q_{2}+2Q_{3}\right)^{2}}{2},$ (61)
there exists the elementary solution of (60)
$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (62)
where
$\psi=\frac{\tau_{R}\left(Q_{1}-Q_{2}\right)}{T^{2}\left(Q_{1}+Q_{2}+2Q_{3}\right)}.$ (63)
In the Appendix A we’ll show how one arrives to the solutions (57) and (62).
Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, E. Konig, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems and Workshop on Localization and Ergodicity, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A The ODE which doesn’t contain explicitly the independent variable
Equation (49) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $Q$ as the new independent variable and
${\cal I}=\frac{dQ}{d\tau}$ (64)
as the new dependent variable. In the new variables (49) takes the form of Abel equation of the second kind polyanin
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=F(Q).$ (65)
As it follows from (39a), the boundary conditions for (65) are
$\lim_{Q\to Q_{1},Q_{2}}{\cal I}(Q)=0.$ (66)
Since (65) is the first order differential equation, the problem is overdetermined. However, for the problems considered explicitly in Section V.1.1, not only will we be able to find the solutions of (65) satisfying the boundary conditions (66) - the solutions will turn out to be simple polynomial (fractional polynomial) functions.
A.1 Integration in terms of elementary functions
For $Q(V)$ given by (59), Eq. (65) can be presented as
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}).$ (67)
The boundary conditions (66) and the r.h.s. of (67) prompt the idea kogan4 ; kogan5 to look for ${\cal I}(Q)$ as the polynomial function proportional to $\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$. Because the r.h.s. of (67) is the third order polynomial of $V$, the multiplying factor should be a constant
${\cal I}(Q)=\frac{\psi}{Q_{1}-Q_{2}}\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$ (68)
(the normalization was chosen for later convenience).
Substituting (68) into (67) we obtain
$\frac{T^{2}\psi^{2}}{\left(Q_{1}-Q_{2}\right)^{2}}\left(2Q-Q_{1}-Q_{2}\right)+\frac{\tau_{R}\psi}{Q_{1}-Q_{2}}=\gamma\left(Q+Q_{3}\right).$
Hence the condition on the parameters of the problem, for (68) to be the solution of (67) is given by (61), and the constant multiplier in (68) should is given by (63). Substituting (68) into Eq. (64) and integrating we obtain the solution (62).
For $Q(V)$ given by (53) we obtain, instead of (67),
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\beta(Q-Q_{1})(Q-Q_{2}).$ (70)
If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial equal to zero at $Q=Q_{1,2}$. So let us split in two one multiplier in the solution (68)
$\frac{Q-Q_{1}}{Q_{1}-Q_{2}}=\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}+1\right].$
This prompts us the idea to look up for the solution of (70) in the form kogan4 ; kogan5
${\cal I}(Q)=2\chi\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left(Q-Q_{2}\right).$ (72)
Substituting the ansatz into (70) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (56) and $\chi$ is given by (58). Substituting thus obtained ${\cal I}(Q)$ into (64) and integrating we obtain (57)
A.2 Solitary wave solution of a generalized KdV equation
To additionally illustrate the method of integration used in this Appendix, let us recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
$\phi_{t}+(n+1)(n+2)\phi^{n}\phi_{x}+\phi_{xxx}=0.$ (73)
For a travelling wave solution
$\phi(t,x)=\phi(\tau),\hskip 28.45274pt\tau=t-x/U,$ (74)
the equation in partial derivatives (73) is reduced to the ODE
$U^{3}\frac{d\phi}{d\tau}-U^{2}(n+1)(n+2)\phi^{n}\frac{d\phi}{d\tau}-\frac{d^{3}\phi}{d\tau^{3}}=0.$ (75)
Equation (75) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $\phi$ as the new independent variable and
$E=\frac{d\phi}{d\tau}$ (76)
as the new dependent variable. As the result, we can present the nonlinear third order ODE (75) as the linear second order ODE for $E^{2}$
$\frac{d^{2}E^{2}}{d\phi^{2}}=2U^{3}-2(n+1)(n+2)U^{2}\phi^{n}.$ (77)
A solitary wave corresponds to the boundary conditions
$\lim_{\tau\to\pm\infty}\phi(\tau)=0,$ (78)
which in the new variables are kogan5
$E^{2}(0)=0,\hskip 28.45274pt\left.\frac{dE^{2}}{d\phi}\right|_{\phi=0}=0.$ (79)
Solving (77) with these boundary conditions we obtain
$E^{2}=U^{3}\phi^{2}-2U^{2}\phi^{n+2}.$ (80)
Substituting thus obtained $E(\phi)$ into (76) and integrating we obtain the well known result drazin
$\phi^{n}=\frac{U}{2}\operatorname{sech}^{2}\left[\frac{n\sqrt{U}}{2}\left(x-Ut\right)\right].$ (81)
Appendix B Shock profile for the fully nonlinear transmission line
For the fully nonlinear transmission line it is convenient to present the problem in the form slightly different from that used everywhere above. We’ll consider the shock speed $U$ and the conditions before the shock as given. Thus the boundary conditions are
$\lim_{\tau\to-\infty}Q(\tau)=Q_{1};\hskip 14.22636pt\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1}$ (82)
and we can concretize (38) to
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=U(Q-Q_{1})-I+I(\Phi_{1}),$ (83a)
$R_{C}\frac{dQ}{d\tau}$ $=U(\Phi-\Phi_{1})-V+V(Q_{1}).$ (83b)
Equation (83) is simplified when there is a leading mechanism of dissipation, in which cases we should substitute into (83) either $R_{C}=0$ or $R_{L}=\infty$. In the first case we equate the r.h.s. of (83b) to zero and obtain
$V=U(\Phi-\Phi_{1})+V(Q_{1}).$ (84)
Substituting $V$ into (3a), solving the equation with respect to $Q$ and then substituting thus obtained $Q(\Phi)$ into (83a) we obtain the equation which can be easily integrated in quadratures. In the second case we equate the r.h.s. of (83a) to zero and obtain
$I=U(Q-Q_{1})+(\Phi_{1}).$ (85)
Substituting $I$ into (3b), solving the equation with respect to $\Phi$ and then substituting thus obtained $\Phi(Q)$ into (83b) we obtain the equation which can be integrated in quadratures. Thus we obtained Eqs. (48a) and (48b) of Ref. kogan3 .
If both $R_{C}>0$ and $R_{L}<\infty$, from (83) we may obtain the equation
$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U(\Phi-\Phi_{1})-V(Q)+V(Q_{1})}{U(Q-Q_{1})-I(\Phi)+I(\Phi_{1})}.$ (86)
If we manage to solve (86), we can substitute the solution into one of the equations (83) and integrate the resulting equation in quadratures to obtain the profile of the shock wave. It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs. (54) and (60) describing weak shocks in half-nonlinear transmission line (i.e. Eqs. (56) and (61)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$. Further analysis of (86) we postpone until later.